OpenStudy (anonymous):

Find a ⋅ b. a = 8i + 6j, b = 4i + 5j <12, 11> 62 -2 <32, 30>

jimthompson5910 (jim_thompson5910):

multiply the corresponding components, then add

jimthompson5910 (jim_thompson5910):

Similar Example a = 7i + 9j, b = 2i + 3j a dot b = 7*2 + 9*3 a dot b = 14 + 27 a dot b = 41 This is NOT your answer as it's just an example.

OpenStudy (anonymous):

is it 62

OpenStudy (anonymous):

thanks for showing how to do this